can you please help with this question...
Joseph is 4 years older than twice his sisters age. In 4 years time he will be 11 years older than here. How old is Joseph now

Answer by venugopalramana(3286)

Joseph is 4 years older than twice his sisters age. In 4 years time he will be 11 years older than her. How old is Joseph now
let joseph's sister's age now be x yrs.
twice her age =2x
4 years older than her would be = 2x+4
this is joseph's age now as per given data.
so joseph's age now =2x+4
after 4 years joseph's sister's age = x+4
11years older to this would be = x+4+11=x+15.......(1)
but after 4 years joseph's age = 2x+4+4=2x+8.......(2)
hence (1)=(2) OR x+15=2x+8
2x+8=x+15
2x-x=15-8=7
x=7....hence joseph's sister's age now =x=7 yrs
and so joseph's age now =2x+4=2*7+4=14+4=18 yrs.